A firm has the following short run total costs, where TC is total cost and Q is output:

Q TC

0 $ 100

1 $ 120

2 160

3 210

4 290

5 390

6 510

7 650

8 810

(a) What is total fixed cost equal to?

(b) What is total variable cost equal to at Q = 2?

(c) What is average total cost equal to at Q = 5?

(d) What is marginal cost equal to at Q =5 ?

(e) At Q =5, is the firm operating under increasing or diminishing returns, and why?
Answer #1

ANSWER:

1) Total fixed cost is equal to the total cost when output is zero as there is no variable cost involved when output is zero and so total fixed cost is $100.

2) Total variable cost at q = 2 will be total cost - total fixed cost.

total variable cost = 160 - 100 = 60

so the total variable cost is $60.

3) Average total cost = total cost / output = 390 / 5 = 78

so the average total cost is $78 (at q = 5)

4) Marginal cost = change in total cost / change in output

marginal cost = ( total cost when output is 5 - total cost when output is 4) / ( output is 5 - output is 4)

marginal cost = (390 - 290) / (5 - 4) = 100 / 1 = 100

so the marginal cost is $100